# LumiMAS: A Comprehensive Framework for Real-Time Monitoring and Enhanced Observability in Multi-Agent Systems

### v1

@@ Abstract
The incorporation of large language models in multi-agent systems (MASs) has the potential to significantly improve our ability to autonomously solve complex problems. However, such systems introduce unique challenges in monitoring, interpreting, and detecting system failures. Most existing MAS observability frameworks focus on analyzing each individual agent separately, overlooking failures associated with the entire MAS. To bridge this gap, we propose LumiMAS, a novel MAS observability framework that incorporates advanced analytics and monitoring techniques. The proposed framework consists of three key components: a monitoring and logging layer, anomaly detection layer, and anomaly explanation layer. LumiMAS’s first layer monitors MAS executions, creating detailed logs of the agents’ activity. These logs serve as input to the anomaly detection layer, which detects anomalies across the MAS workflow in real time. Then, the anomaly explanation layer performs classification and root cause analysis (RCA) of the detected anomalies. LumiMAS was evaluated on seven different MAS applications, implemented using two popular MAS platforms, and a diverse set of possible failures. The applications include two novel failure-tailored applications that illustrate the effects of a hallucination or bias on the MAS. The evaluation results demonstrate LumiMAS’s effectiveness in failure detection, classification, and RCA.
1 Introduction
The incorporation of large language models (LLMs) in multi-agent systems (MASs) is rapidly reshaping intelligent systems by enabling dynamic context-aware interactions across real-world applications (Guo et al. 2024). For instance, MetaGPT (Hong et al. 2023) demonstrates how LLM-based MASs can coordinate effectively to solve complex coding tasks. Another example is PMC (Zhang et al. 2024a), which assigns constraint-specific planning agents that solve real-world problems such as travel planning.
@@ 1 Introduction
However, the incorporation of LLMs makes MASs susceptible to the vulnerabilities inherent in these models, such as (1) hallucinations (Ji et al. 2023), (2) biases (Navigli, Conia, and Rossini 2023), and (3) adversarial manipulations (Zhang et al. 2023) that threaten the trustworthiness of AI systems. Since errors can propagate through inter-agent interactions, such issues are often amplified in MASs (Stoica et al. 2025). Furthermore, the autonomous and dynamic nature of MASs exposes them to risks highlighted by the OWASP Top 10 for LLM Applications Project (OWASP Foundation 2025), including prompt injections, memory poisoning, and cascading hallucinations. Notably, the 2025 OWASP list underscores the importance of robust monitoring solutions in managing agentic AI risks. Nowadays, most existing monitoring and failure detection approaches analyze each agent separately, utilize LLM that induce latency and computational cost to the system, or focus on a limited set of possible failures, which ultimately harm the relevance of the solution, limiting their effectiveness for real-time, multi-user systems and not aligned with the users’ needs.
To address this gap, we present LumiMAS, a novel agnostic and efficient framework that incorporates advanced analytics and monitoring techniques, comprehensively capturing both system-level features and the semantic nuances of inter-agent interactions. LumiMASdesigned to enhance observability, improves monitoring capabilities, enables real-time failure detection with minimal resource consumption, and supports the identification of a diverse range of system failures, as well as to improve the model’s alignment with the user’s needs. LumiMAS is composed of three key components: (1) a monitoring layer that performs system-wide logging across diverse MAS frameworks (extracting key operational and communication features); (2) an anomaly detection layer which identifies deviations from normal behavior in real-time; and (3) an anomaly explanation layer that consists of anomaly classification and root cause analysis (RCA) LLM-based agents (LMAs) whose aim is to identify the anomaly’s type and source.
LumiMAS was evaluated on seven MAS applications, implemented using two popular MAS platforms, and a diverse set of failures. The evaluation results demonstrate LumiMAS’s abilities to detect failures in real-time (under 0.07 seconds) and with high accuracy and a low average false positive rate. Additionally, we introduce novel failure-tailored applications (FTAs) to reliably benchmark failures, such as hallucinations and bias, bridging the evaluation gap between theoretical research and real-world deployment.
Our contributions can be summarized as follows: (1) To the best of our knowledge, we are the first to propose a novel platform-agnostic observability framework for LLM-based MASs, which enables monitoring and detecting issues in real time, such as adversarial attacks, bias, and hallucinations, that threaten AI systems’ reliability. (2) We propose novel FTAs to evaluate detection and mitigation strategies for common MAS failures. (3) Our extensive evaluation across seven MAS applications using two MAS platforms demonstrates LumiMAS’s effectiveness in anomaly detection, classification, and root cause analysis. (4) A unique dataset of the collected logs from the various applications examined in this work, made available to facilitate further research in this domain.
2 Related Work
@@
Despite the rapid adoption of LLM-based MASs, limited research has been conducted to assess their risks, mitigate failures, and monitor their workflow. However, the increased use of LMAs exposes the system they operate within to new threats; highlighting the need for robust tools to effectively monitor and validate their actions.
2.1 MAS Risk Assessment
@@
To enhance MAS safety evaluation, Ruan et al. (2023) used LLMs to emulate sandbox environments, enabling pre-deployment testing of agents when interacting with new tools. The work was extended by Yuan et al. (2024), who added ambiguous instructions to evaluate whether agents can distinguish safe from unsafe actions. Zhang et al. (2024b) introduced the agent security bench, which evaluates attacks and defenses on LMAs operating in a MAS. While these studies concentrated on assessing the safety of MASs, our approach monitors deployed MASs to improve system observability, capturing a broad range of execution features, and facilitating real-time failure detection.
2.2 Failure Mitigation Strategies
@@
Employing rule‑based mitigation strategies for unsafe actions, AgentMonitor (Naihin et al. 2023), TrustAgent (Hua et al. 2024), and the visibility framework of Chan et al. (2024a), continuously inspect the LMA’s chain‑of‑thought and planned tool calls in real time, to score or block unsafe steps and enforce domain‑specific safety policies. Leveraging the reasoning capabilities of LLMs, Fang, Zhu, and Gurevych (2024) introduced an approach to infer the agent’s intent from its action sequence, which is then verified to align with the user’s instructions. Recently Peigné et al. (2025) highlight how malicious prompts can propagate across agents in MAS, and introduce individual-agent mitigation strategies demonstrating a trade-off between robustness and cooperation. However, unlike these approaches, which primarily focus on individual agents operating in a MAS, LumiMAS addresses critical inter-agent dynamics and system-level failure detection.
2.3 MAS Monitoring
@@
AgentOps (Dong, Lu, and Zhu 2024) lays the theoretical groundwork for observing LMAs by tracing key agent artifacts and their associated data throughout the agent’s life cycle. Chan et al. (2024) proposed a solution for monitoring individual agent inputs and outputs and applying real-time corrections. In addition, a detailed list of commercial observability tools (e.g., LangSmith, LangFuse) can be found in Appendix A.
@@ 3 Threat Model
LLM-based MASs, which incorporate multiple components, are vulnerable to various failures that either originate from inherent limitations of the LLM, such as hallucinations (Ji et al. 2023) and biases (Xu et al. 2025; Navigli, Conia, and Rossini 2023), or result from adversarial interference (Deng et al. 2025; OWASP Foundation 2025).
@@
Adversarial goal: As described by Zhang et al. (2024b), the primary adversarial goal is to manipulate AI agents into performing actions that deviate from their intended behavior, potentially resulting in unsafe, unethical, or harmful outcomes. Adversaries may attempt to cause system failures, degrade overall system performance, extract sensitive information, or perform resource exhaustion.
@@
Threat actors, their capabilities and attack vectors: Similar to Deng et al. (2025) and Zhang et al. (2024b), we consider a diverse set of threat actors that reflect realistic risks in MAS. While all threat actors are assumed to have prior knowledge, which could have been obtained in a prior reconnaissance stage, they typically lack access to the internals of the underlying LLM (i.e., the LLM is treated as a black-box). The adversarial attack can be performed by malicious users which may utilize different attack vectors including direct prompt injection (DPI) and indirect prompt injection (IPI). In a DPI attack, the threat actor manipulates the input prompt to inject malicious instructions, misinformation, or even backdoor commands to compromise the system’s integrity. IPI attacks exploit the agent’s reliance on external tools (e.g., web search), allowing adversaries to plant malicious instructions within these sources and cause unintended actions. The attack may also be executed by external data providers or privileged users who control data sources (e.g., retrieval-augmented generation (RAG) database). In that case, an attacker may perform a memory poisoning (MP) attack, in which the agent’s memory source is contaminated, causing the agent to behave according to the attacker’s intent.
4 Methodology
@@
To address the challenges in monitoring and interpreting MAS behavior, we propose a comprehensive, multi-layered framework to enhance system observability. This methodology comprises three key components: (1) a monitoring and logging layer, (2) an anomaly detection layer, and (3) an anomaly explanation layer consisting of anomaly classification and RCA techniques.
@@
In LumiMAS’s first phase, execution data is systematically collected, aiming to capture diverse low-level operational features and high-dimensional semantic information reflecting agent activities and reasoning. The collected data undergoes preprocessing, transforming it into structured formats suitable for quantitative analysis and enabling understanding of complex system dynamics. LumiMAS’s analytical techniques employed in the anomaly detection and explanation layers utilize this processed data.
@@
Figure 1: High-level architecture of the LumiMAS observability framework incorporated in a given MAS.
4.1 System Logging Layer
@@
A fundamental component of our observability framework is a comprehensive system logging layer. LumiMAS logs several unique predefined events that abstract away implementation-specific behaviors, ensuring compatibility with different MAS platforms. The content of each log encodes system behavior into a chronological sequence of structured events $\{e_{1},e_{2},\dots,e_{n}\}$ where each $e_{i}$ is an event selected from a predefined set of event types $E$. The life cycle of a MAS application is bounded by the events Application-Started and Application-Ended. Within these boundaries, agent-specific execution spans are recorded via Agent-Started(a)(a) and Agent-Finished(a)(a), where $a\in{A}$ denotes an agent identity from the set of all agents A{A}. Throughout the agent’s execution, we log every LLM invocation and tool usage as discrete events (LLM-Call and Tool-Usage) for every iteration, capturing the agent’s behavior. The telemetry data collected for each of these events includes operational features detailing agents’ behavior (e.g., number of tool invocations), resource consumption (e.g., token usage counts), and semantic information derived from agents’ LLM interactions. Each event’s monitored features are provided in Appendix B. Our logging approach is designed to be agnostic to the specific underlying MAS platform (e.g., CrewAI (CrewAI Inc. 2025), LangGraph (LangGraph 2023)).
4.2 Anomaly Detection Layer
@@
The next component of our framework is an anomaly detection layer designed to be lightweight (to facilitate real-time detection) and identify various failures in MAS applications. Our detection method combines engineered features that capture structural and statistical properties of agent behavior and semantics features extracted from the text generated during agent interactions (e.g., LLM outputs) as well as leverages an LSTM-based autoencoder (AE) architecture, which has proven particularly effective for the anomaly detection task (Lee et al. 2024; Wei et al. 2023). Under the assumption that failures within MAS exhibit abnormal behavior in agent execution and semantic communication, we propose three AE-based detection approaches: (1) Execution performance indicator (EPI) detection, which utilizes features extracted from each agent’s execution, such as execution duration and token count (features are elaborated in Appendix B ). (2) Semantic detection, which utilizes encoded representations of outputs from LLM interactions (e.g., the agent’s reasoning process). (3) Combined latent-space detection, which employs a linear AE trained on the concatenated latent representations from the first two approaches. Figure 2 presents our anomaly detection architecture overview. The AE models are trained to reconstruct features from logs collected using our system logging layer. At inference, the features extracted from the execution logs are propagated through the model, which returns reconstructed representations of the features. The reconstruction error between the model’s input and output is computed; a high reconstruction error indicates potential anomalies. The sequential nature of our anomaly detection approach is used to capture the whole system’s normal patterns and flow, which helps detect deviations from those patterns. By combining these different representations, our method can detect a wide range of anomalies, without being limited to a specific application or failure type.
Feature Extraction.
@@
Figure 2: anomaly detection architecture overview: (1) EPI, (2) Semantic, and (3) Combined detection AEs.
Model Training.

### v2

# arXiv:2508.12412v2[cs.CR] 27 Feb 2026

@@ Abstract.
The incorporation of LLMs in multi-agent systems (MASs) has the potential to significantly improve our ability to autonomously solve complex problems. However, such systems introduce unique challenges in monitoring, interpreting, and detecting system failures. Most existing MAS observability frameworks focus on analyzing each individual agent separately, overlooking failures associated with the entire MAS. To bridge this gap, we propose LumiMAS, a novel MAS observability framework that incorporates advanced analytics and monitoring techniques. The proposed framework consists of three key components: a monitoring and logging layer, anomaly detection layer, and anomaly explanation layer. LumiMAS’s first layer monitors MAS executions, creating detailed logs of the agents’ activity. These logs serve as input to the anomaly detection layer, which detects anomalies across the MAS workflow in real time. Then, the anomaly explanation layer performs classification and root cause analysis (RCA) of the detected anomalies. LumiMAS was evaluated on seven different MAS applications, implemented using two popular MAS platforms, and a diverse set of possible failures. The applications include two novel failure-tailored applications that illustrate the effects of a hallucination or bias on the MAS. The evaluation results demonstrate LumiMAS’s effectiveness in failure detection, classification, and RCA.
Key words and phrases:
AI Agents; AI Security; Anomaly Detection; Observability
doi:
PNWL9707
1. Introduction
@@
The incorporation of large language models (LLMs) in multi-agent systems (MASs) is rapidly reshaping intelligent systems by enabling dynamic context-aware interactions across real-world applications guo2024large. For instance, MetaGPT hong2023metagpt demonstrates how LLM-based MASs can coordinate effectively to solve complex coding tasks. Another example is PMC zhang2025planning, which assigns constraint-specific planning agents that solve real-world problems such as travel planning. However, the incorporation of LLMs makes MASs susceptible to the vulnerabilities inherent in these models, such as (1) hallucinations ji2023survey, (2) biases navigli2023biases, and (3) adversarial manipulations greshake2023not that threaten the trustworthiness of AI systems. Since errors can propagate through inter-agent interactions, such issues are often amplified in MASs cemri2025multi. Furthermore, the autonomous and dynamic nature of MASs exposes them to risks highlighted by the OWASP Top 10 for LLM Applications Project owaspLLMTop10, including prompt injections, memory poisoning, and cascading hallucinations. Notably, the 2025 OWASP list underscores the importance of robust monitoring solutions and srikumar2025agents srikumar2025agents highlight the need for real-time failure detection solutions in managing agentic AI risks. Nowadays, most existing monitoring and failure detection approaches analyze each agent separately, utilize LLM that induce latency and computational cost to the system, or focus on a limited set of possible failures, which ultimately harm the relevance of the solution, limiting their effectiveness for real-time, multi-user systems, and are not aligned with the users’ needs. To address this gap, we present LumiMAS, a novel agnostic and efficient framework that incorporates advanced analytics and monitoring techniques, comprehensively capturing both system-level features and the semantic nuances of inter-agent interactions. LumiMAS utilize existing techniques with engineered semantic and operational features to effectively address the open problem of real-time MAS failure detection. It is designed to enhance observability, improve monitoring capabilities, enable real-time failure detection with minimal resource consumption, and supports the identification of a diverse range of system failures, as well as to improve the model’s alignment with the user’s needs. LumiMAS is composed of three key components: (1) a monitoring layer that performs system-wide logging across diverse MAS frameworks (extracting key operational and communication features); (2) an anomaly detection layer which identifies deviations from normal behavior in real-time; and (3) an anomaly explanation layer that consists of anomaly classification and root cause analysis (RCA) LLM-based agents (LMAs) whose aim is to identify the anomaly’s type and source.
2. Related Work
@@
Despite the rapid adoption of LLM-based MASs, limited research has been conducted to assess their risks, mitigate failures, and monitor their workflow. However, the increased use of LMAs exposes the system they operate within to new threats, highlighting the need for robust tools to monitor and validate their actions effectively.
2.1. MAS Risk Assessment
@@
To enhance MAS safety evaluation, ruan2023identifying ruan2023identifying used LLMs to emulate sandbox environments, enabling pre-deployment testing of agents when interacting with new tools. The work was extended by yuan2024r yuan2024r, who added ambiguous instructions to evaluate whether agents can distinguish safe from unsafe actions. zhang2024agent zhang2024agent introduced the agent security bench, which evaluates attacks and defenses on LMAs operating in a MAS. While these studies concentrated on assessing the safety of MASs, our approach monitors deployed MASs to improve system observability, capturing a broad range of execution features, and facilitating real-time failure detection.
2.2. Failure Mitigation Strategies
@@
Using rule‑based mitigation strategies for unsafe actions, AgentMonitor naihin2023testing, TrustAgent hua2024trustagent, AgenTRIM betser2026agentrim, and the visibility framework of chan2024visibility chan2024visibility, continuously inspect the LMA’s chain‑of‑thought and planned tool calls in real time, to score or block unsafe steps and enforce domain‑specific safety policies. Leveraging the reasoning capabilities of LLMs, fang2407inferact fang2407inferact introduced an approach to infer the agent’s intent from its action sequence, which is then verified to align with the user’s instructions. Recently peigne2025multi peigne2025multi highlight how malicious prompts can propagate across agents in MAS, and introduce individual-agent mitigation strategies demonstrating a trade-off between robustness and cooperation. However, unlike these approaches, which primarily focus on individual agents operating in a MAS, LumiMAS addresses critical inter-agent dynamics and system-level failure detection.
2.3. MAS Monitoring
@@
AgentOps dong2024agentops lays the theoretical groundwork for observing LMAs by tracing key agent artifacts and their associated data throughout the agent’s life cycle. chan2024agentmonitor chan2024agentmonitor proposed a solution for monitoring individual agent inputs and outputs and applying real-time corrections. In addition, a detailed list of commercial observability tools (e.g., LangSmith, LangFuse) can be found in Appendix A.
@@ 3. Threat Model
LLM-based MASs, which incorporate multiple components, are vulnerable to various failures that either originate from inherent limitations of the LLM, such as hallucinations ji2023survey and biases xu2025mitigating; navigli2023biases, or result from adversarial interference deng2025ai; owaspLLMTop10.
Adversarial goal
@@
As described by zhang2024agent zhang2024agent, the primary adversarial goal is to manipulate AI agents into performing actions that deviate from their intended behavior, potentially resulting in unsafe, unethical, or harmful outcomes. Adversaries may attempt to cause system failures, degrade overall system performance, extract sensitive information, or perform resource exhaustion.
Threat actors, their capabilities and attack vectors
@@
Similar to deng2025ai deng2025ai and zhang2024agent zhang2024agent, we consider a diverse set of threat actors that reflect realistic risks in MAS. While all threat actors are assumed to have prior knowledge, which could have been obtained in a prior reconnaissance stage, they typically lack access to the internals of the underlying LLM (i.e., the LLM is treated as a black-box). The adversarial attack can be performed by malicious users who may utilize different attack vectors, including direct prompt injection (DPI) and indirect prompt injection (IPI). In a DPI attack, the threat actor manipulates the input prompt to inject malicious instructions, which can lead to resource exhaustion, propagation of misinformation, or insertion of backdoor commands to compromise the system’s integrity. IPI attacks exploit the agent’s reliance on external tools (e.g., web search), allowing adversaries to plant malicious instructions within these sources and cause unintended actions. The attack may also be executed by external data providers or privileged users who control data sources (e.g., retrieval-augmented generation (RAG) database). In that case, an attacker may perform a memory poisoning (MP) attack, in which the agent’s memory source is contaminated, causing the agent to behave according to the attacker’s intent.
3.1. Threat Taxonomy
According to cemri2025multi cemri2025multi, who introduced the MAST taxonomy, multi-agent failures can be classified into three categories: Specification Issues, Inter-Agent Misalignment, and Verification or Termination Deficiencies. When viewed through this lens, each failure type aligns with specific categories and failure modes defined in the MAST taxonomy. Our examined failures are mostly aligned with the first two categories, with limited focus on the third (Task Verification). Under specification issues, the following modes are most relevant: FM-1.1 Disobey Task Specification – failure to adhere to task constraints or requirements (DPI exhaustion, IPI); FM-1.2 Disobey Role Specification – deviation from assigned responsibilities or confusion between roles (DPI backdoor); FM-1.3 Step Repetition – redundant task execution or looping (DPI exhaustion, IPI); and FM-1.4 Loss of Conversation History – truncation or overwriting of prior context (DPI misinformation, MP, hallucination). Under Inter-Agent Misalignment, the relevant modes include FM-2.1 Conversation Reset – dialogue restart with context loss (hallucination, IPI); FM-2.3 Task Derailment – deviation from the intended goal (DPI, IPI, MP, hallucination); FM-2.4 Information Withholding – failure to share critical data across agents (IPI, DPI); FM-2.5 Ignored Other Agent’s Input – disregard of peer contributions (IPI, DPI, MP); and FM-2.6 Reasoning–Action Mismatch – inconsistency between reasoning and behavior (MP, hallucination). In our work, we primarily target failures in specification integrity and misalignment, while explicit verification-oriented mechanisms remain a promising area for future extension.
4. Methodology
@@
To address the challenges in monitoring and interpreting MAS behavior and addressing the open problem of real-time failure detection, we propose a comprehensive, multi-layered framework to enhance system observability. This methodology comprises three key components: (1) a monitoring and logging layer, (2) an anomaly detection layer, and (3) an anomaly explanation layer consisting of anomaly classification and RCA techniques.
@@
In LumiMAS’s first phase, execution data is systematically collected, aiming to capture diverse low-level operational features and high-dimensional semantic information reflecting agent activities and reasoning. The collected data undergoes preprocessing, transforming it into structured formats suitable for quantitative analysis and enabling an understanding of complex system dynamics. LumiMAS’s analytical techniques employed in the anomaly detection and explanation layers utilize this processed data.
@@
Figure 1. High-level architecture of the LumiMAS observability framework incorporated in a given MAS
[…]
High-level architecture of the LumiMAS observability framework integrated into a multi-agent system (MAS). System input is processed by the MAS, which consists of multiple interacting agents. The framework monitors the MAS through system logging and performs anomaly detection and anomaly explanation.
4.1. System Logging Layer
@@
A fundamental component of our observability framework is a comprehensive system logging layer. LumiMAS logs several unique predefined events that abstract away implementation-specific behaviors, ensuring compatibility with different MAS platforms. The content of each log encodes system behavior into a chronological sequence of structured events $\{e_{1},e_{2},\dots,e_{n}\}$ where each $e_{i}$ is an event selected from a predefined set of event types $E$. The life cycle of a MAS application is bounded by the events Application-Started and Application-Ended. Within these boundaries, agent-specific execution spans are recorded via Agent-Started(a)(a) and
Agent-Finished(a)(a), where $a\in{A}$ denotes an agent identity from the set of all agents A{A}. Throughout the agent’s execution, we log every LLM invocation and tool usage as discrete events (LLM-Call and Tool-Usage) for every iteration, capturing the agent’s behavior. The telemetry data collected for each of these events includes operational features detailing agents’ behavior (e.g., number of tool invocations), resource consumption (e.g., token usage counts), and semantic information derived from agents’ LLM interactions. Each event’s monitored features are provided in Appendix B. Our logging approach is designed to be agnostic to the specific underlying MAS platform (e.g., CrewAI crewai, LangGraph LangGraphMisc).
4.2. Anomaly Detection Layer
@@
The next component of our framework is an anomaly detection layer designed to be lightweight (to facilitate real-time detection) and identify various failures in MAS applications. Our detection method combines engineered features that capture structural and statistical properties of agent behavior and semantics features extracted from the text generated during agent interactions (e.g., LLM outputs). As our method designed to handle long-term dependencies, inter-agent communication patterns, and variable-length logs in a dynamic multi-agent environment. To maintain low latency under these conditions, we employ an LSTM-based autoencoder (AE) architecture, which has proven particularly effective for the anomaly detection task lee2024lstm; wei2023lstm. Under the assumption that failures within MAS exhibit abnormal behavior in agent execution and semantic communication, we propose three AE-based detection approaches: (1) Execution performance indicator (EPI) detection, which utilizes features extracted from each agent’s execution, such as execution duration and token count (features are elaborated in Appendix B). (2) Semantic detection, which utilizes encoded representations of outputs from LLM interactions (e.g., the agent’s reasoning process). (3) Combined latent-space detection, which employs a linear AE trained on the concatenated latent representations from the first two approaches. Figure 2 presents our anomaly detection architecture overview. The AE models are trained to reconstruct features from logs collected using our system logging layer. At inference, the features extracted from the execution logs are propagated through the model, which returns reconstructed representations of the features. The reconstruction error between the model’s input and output is computed; a high reconstruction error indicates potential anomalies. The sequential nature of our anomaly detection approach is used to capture the whole system’s normal patterns and flow, which helps detect deviations from those patterns. By combining these different representations, our method can detect a wide range of anomalies, without being limited to a specific application or failure type.
4.2.1. Feature Extraction
@@
Figure 2. Overview of the anomaly detection architecture showing (1) the EPI-based autoencoder, (2) the semantic-based autoencoder, and (3) a combined detector that integrates both approaches
[…]
Overview of the anomaly detection architecture with three variants: an EPI-based autoencoder, a semantic-based autoencoder, and a combined detector that integrates both approaches for anomaly detection. The figure illustrates the flow of information through each variant and highlights the shared and distinct components of the three designs.
4.2.2. Model Training